An electron moves in uniform magnetic field with a speed of 5·106 m/s along the x axis. The magnetic field strength is 0.050 T, directed at an angle of 45° to the x axis and lying in the xy plane. Calculate the magnetic force 𝐹 on and acceleration 𝑎 of the electron.


1
Expert's answer
2021-05-27T08:29:40-0400

(a) We can find the magnetic force on the electron as follows:


"F=qvBsin\theta,""F=1.6\cdot10^{-19}\ C\cdot5\cdot10^6\ \dfrac{m}{s}\cdot0.05\ T\cdot sin45^{\circ}=2.83\cdot10^{-14}\ N."

(b) We can find the acceleration of the electron from Newton's Second Law of Motion:


"F=ma,""a=\dfrac{F}{m}=\dfrac{2.83\cdot10^{-14}\ N}{9.11\cdot10^{-31}\ kg}=3.1\cdot10^{16}\ \dfrac{m}{s^2}."